OpenStudy (anonymous):

which equation represents the line passing through the points (1,1) (-2,13)? PLEASE HELP!

OpenStudy (anonymous):

did you find the slope?

OpenStudy (anonymous):

no I haven't but I know how to I jusr don't know the formula by heart

OpenStudy (anonymous):

it is "the change in y over the change in x" or as a formula \[\frac{y_2-y_1}{x_2-x_1}\] replace \(y_2\) by \(13\), \(y_1\) by \(1\), \(x_2\) by \(-2\) and \(x_1\) by \(1\)

OpenStudy (anonymous):

so -4?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

-4 is the slope

OpenStudy (anonymous):

ya I know but I don't know what to do next..thank you for replying

OpenStudy (anonymous):

you have to replace one of the points in the equation.. The equation have the next form: \[\frac{ y _{2}-y _{1} }{x _{2}-x _{1}}=m \] \[\frac{ y _{2}-1 }{x _{2}-1}=-4\]

OpenStudy (anonymous):

clearing .. the equation of this is y=-4x+5

OpenStudy (anonymous):

thank for your help!

OpenStudy (anonymous):

that's not answer on my choices

OpenStudy (anonymous):

which are your choices?

OpenStudy (anonymous):

4x-y=5 x-4y+5 x+4y=5 4x+y=5

OpenStudy (anonymous):

you have to move "4x" to the other member of the equation y+4x=5 . The last choice

OpenStudy (anonymous):

ohh ok...thank you!!